### v1

# On Multilingual Encoder Language Model Compression for Low-Resource Languages

Abstract
In this paper, we combine two-step knowledge distillation, structured pruning, truncation, and vocabulary trimming for extremely compressing multilingual encoder-only language models for low-resource languages. Our novel approach systematically combines existing techniques and takes them to the extreme, reducing layer depth, feed-forward hidden size, and intermediate layer embedding size to create significantly smaller monolingual models while retaining essential language-specific knowledge. We achieve compression rates of up to 92% with only a marginal performance drop of 2–10% in four downstream tasks, including sentiment analysis, topic classification, named entity recognition, and part-of-speech tagging, across three low-resource languages. Notably, the performance degradation correlates with the amount of language-specific data in the teacher model, with larger datasets resulting in smaller performance losses. Additionally, we conduct extensive ablation studies to identify best practices for multilingual model compression using these techniques.
1 Introduction
Small multilingual encoder language models (LMs), such as mBERT Devlin et al. (2019), XLM-R Conneau et al. (2020), and Glot-500m Imani et al. (2023), have demonstrated strong performance across a diverse range of low-resource languages Hu et al. (2020); Asai et al. (2024), often outperforming large-scale proprietary models on various sequential tasks Adelani et al. (2024); Gurgurov et al. (2025). However, even these relatively compact multilingual models may still be excessively large for use in individual languages due to redundant capacity and expensive inference Singh and Lefever (2022); Cruz (2025).
To address this, we propose a novel combination of model compression approaches for transforming multilingual encoder-only models into maximally small, efficient, language-specific alternatives while retaining competitive performance. Our methodology integrates knowledge distillation Hinton et al. (2015), structured pruning Kim and Hassan (2020); Hou et al. (2020), weight truncation, and vocabulary trimming Abdaoui et al. (2020); Ushio et al. (2023) to systematically reduce model size by compressing the depth (number of layers), feed-forward intermediate width, hidden size, and the tokenizer vocabulary. Our experiments demonstrate that this pipeline achieves compression rates of up to 92%, with only a minor performance drop of 2–10% on downstream tasks such as sentiment analysis, topic classification, named entity recognition, and part-of-speech tagging. Notably, the extent of degradation depends more on the strength of the teacher model than on the compression itself.
Figure 1: Overview of our multilingual model compression methodology. We use (1) knowledge distillation to reduce layers, (2) structured pruning to eliminate redundant feed-forward network width, and (3) hidden size reduction and another round of knowledge distillation from the previous student model. Finally, (4) vocabulary trimming is applied to retain language-specific tokens.
Beyond compression, we investigate the impact of using multilingual versus monolingual teacher models, evaluate different initialization strategies for knowledge distillation, and analyze additional compression variables. Our findings contribute to the development of highly efficient environmentally friendly models Strubell et al. (2020) for low-resource languages and explores how strongly models can be compressed.
2 Methodology
In this section, we present our multilingual model compression strategy, illustrated in Figure 1. Our approach combines several existing compression techniques in a novel way that, to the best of our knowledge, has not been explored in this combination within the multilingual context.
2.1 Layer Reduction via Knowledge Distillation
We reduce the number of transformer layers in the teacher model by half to obtain an initial compact student model Sanh et al. (2020). The student is initialized with the layers of the teacher and trained using a combination of Masked Language Modeling (MLM) Devlin et al. (2019) and Mean Squared Error (MSE) loss for knowledge distillation Hinton et al. (2015). Both losses are weighted equally (0.5). The teacher is a multilingual encoder fine-tuned on the target language (see Section 4).
2.2 Width Reduction via Structured Pruning
We apply structured pruning Kim and Hassan (2020) to reduce the intermediate size of the feed-forward layers from 3072 to 2048. Neuron importance is estimated using first-order gradient information accumulated from forward and backward passes over MLM validation data. At each layer, neurons are ranked by their absolute gradient values, and the least important ones are removed based on a target pruning ratio.
2.3 Hidden Size Compression with Secondary Knowledge Distillation
We compress the hidden embedding dimension from 768 to either 312, 456, or 564 via truncation, retaining the first $k$ dimensions. A second round of knowledge distillation is then performed, using the width-reduced model from the previous step as the new teacher Wang et al. (2023).
2.4 Vocabulary Reduction
We reduce the vocabulary size by selecting the top 40,000 most frequent tokens from a target-language corpus, along with their corresponding embeddings Ushio et al. (2023). This ensures that the resulting model retains only language-specific tokens.
3 Experiments
Below, we describe the datasets, languages, tasks, and baseline systems used in our evaluation.
3.1 Knowledge Distillation Data
We use GlotCC Kargaran et al. (2025), a large-scale multilingual corpus derived mainly from CommonCrawl Wenzek et al. (2020), as the primary dataset for both stages of knowledge distillation. Data distributions for the selected languages are reported in Appendix F. We use GlotCC for training, and the FLORES-200 development set Team et al. (2022) for validation during training.
3.2 Languages and Tasks
We evaluate our models on four tasks: Topic Classification (TC), Sentiment Analysis (SA), Named Entity Recognition (NER), and Part-of-Speech Tagging (POS), covering three low-resource languages—Maltese, Slovak, and Swahili Joshi et al. (2020). For TC, we use the 7-class SIB-200 dataset Adelani et al. (2024), and for SA, we compile binary sentiment datasets from multiple sources Dingli and Sant (2016); Cortis and Davis (2019); Pecar et al. (2019); Muhammad et al. (2023a, b). For NER, we use WikiANN Pan et al. (2017), and for POS, we use Universal Dependencies v2.15 de Marneffe et al. (2021) and MasakhaPOS Dione et al. (2023). For all tasks, we train Sequential Bottleneck task adapters Pfeiffer et al. (2020) with fixed hyperparameters (see Appendix H). Performance is measured using macro-averaged F1 Sokolova et al. (2006) for TC and SA, and "seqeval" F1 Nakayama (2018) for NER and POS.
3.3 Models and Baselines
We compress two encoder-only multilingual models–mBERT Devlin et al. (2019) and XLM-R-base Conneau et al. (2020)–adapted to target languages through fine-tuning on language-specific data, and compare the reduced models to two baselines: (1) the original, non-adapted models, and (2) language-adapted versions. In both cases, we train an identical task adapter using the same task-specific datasets as for the compressed models.
4 Findings
Our key findings are outlined below.
4.1 Distillation
Distilling knowledge from a multilingual teacher into a monolingual student model is less effective than using a target-language adapted teacher, as evidenced by the differences in validation accuracies shown in Figure 2. This discrepancy possibly stems from the multilingual teacher’s broad cross-lingual representations, which are not directly aligned with the requirements of a monolingual student. In contrast, monolingual teachers provide more targeted, language-specific representations, resulting in better student performance.
(a) Maltese (mBERT)
(b) Maltese (XLM-R)
(c) Slovak (mBERT)
(d) Slovak (XLM-R)
(e) Swahili (mBERT)
(f) Swahili (XLM-R)
Figure 2: First-step KD validation accuracies for mBERT and XLM-R with models initialized using the last $k$ layers. mBERT- and XLM-R-mt, sk, sw refer to models adapted to the target language; distilled denotes models trained with distillation loss, while student refers to identically trained models without distillation loss. The best accuracy is in all cases achieved when distilling from a target-language adapted model.
Distillation loss: We compare KL divergence and MSE as distillation loss functions, and observe that MSE leads to better and faster convergence (Appendix A), in line with prior work Kim et al. (2021); Nityasya et al. (2022).
4.2 Weight Initialization
Weight initialization plays a crucial role in training the student model, with knowledge distillation providing only a marginal additional performance improvement (Figure 2). This partly aligns with the findings of Wibowo et al. (2024), who explored distilling multilingual abilities for multilingual tasks, whereas our focus is on monolingual distillation. Training a student-sized model initialized with teacher weights, but without knowledge distillation, results in only a slight performance drop compared to a fully distilled model.
Initialization Strategies: Among various initialization strategies, initializing the student with the last $k$ layers for mBERT and every other layer (stride) for XLM-R consistently outperforms alternatives such as using the first $k$ layers and combining first and last layers (Appendix B). Random initialization performs significantly worse, emphasizing the importance of weight reuse Sun et al. (2019); Singh and Lefever (2022).
4.3 Pruning and Truncation
Distilled models can be compressed further using structured pruning, hidden size reduction, and vocabulary trimming, while maintaining competitive performance.
Intermediate size reduction: Reducing the intermediate size of feed-forward layers from 3072 to 2048 via structured pruning results in negligible performance loss (Table 1). However, more aggressive reductions degrade quality significantly, making 2048 a practical lower bound. We do not prune attention heads, as removing even a minimal number (e.g., three) causes severe degradation.
Hidden size reduction: We reduce the hidden embedding size to 564, 456, and 312, [^1] The hidden size must be divisible by the number of attention heads. truncating it to the first $k$ dimensions. Training is performed under the supervision of the student from the previous stage. We find that using the original teacher leads to worse results, possibly due to the bigger knowledge gap Wang et al. (2023). We also tested SVD-based dimensionality reduction but found truncation to be more effective (see Appendix C).
Model Configuration Red. Size XLM-R-base Avg↑
mt sk sw
TC↑ SA↑ NER↑ POS↑ TC↑ SA↑ NER↑ POS↑ TC↑ SA↑ NER↑ POS↑
Base 0% 279M 68.12 56.04 54.29 89.94 88.11 95.55 91.09 97.31 78.44 81.46 84.64 89.38 81.20
Base-[mt, sk, sw] 0% 279M 84.96 76.20 69.18 95.41 86.23 94.79 91.01 97.09 87.52 84.08 82.67 89.22 86.53
KD layer red. ×2 -15% 236M 84.05 77.20 63.47 94.25 86.33 92.88 90.12 96.26 82.89 81.29 82.85 89.22 85.07
inter. layer red. → 2048 -20% 226M 84.74 78.63 60.12 94.16 86.14 93.35 89.98 96.05 82.39 82.70 83.62 89.54 85.12
* KD hid. size red. → 564 -40% 163M 83.43 74.86 52.97 93.73 84.89 92.74 89.09 96.78 85.84 80.96 80.80 89.42 83.79
vocab. red. → 40k -85% 45M 84.11 72.39 60.92 93.00 85.29 92.92 89.25 96.42 85.71 80.94 81.96 89.12 84.34
* KD hid. size red. → 456 -53% 131M 78.48 69.91 62.54 92.72 85.99 92.99 88.30 96.28 83.06 79.25 80.69 88.94 83.26
vocab. red. → 40k -87% 35M 78.47 70.68 63.25 92.51 86.10 92.91 88.35 96.29 82.49 79.02 80.19 89.00 83.27
* KD hid. size red. → 312 -68% 89M 66.94 70.05 35.73 87.58 83.98 90.93 88.02 95.51 76.37 80.11 80.66 88.27 78.68
vocab. red. → 40k -92% 23M 67.22 71.35 37.06 87.49 84.00 90.52 88.20 95.56 77.95 80.47 79.18 88.01 78.92
Table 1: Performance comparison before and after model reduction for XLM-R-base across Maltese (mt), Slovak (sk), and Swahili (sw) for TC, SA, NER, and POS. Red. shows the parameter reduction compared to the base model. All scores are averaged over three runs with different random seeds. The results for mBERT are in Appendix J.
Vocabulary trimming: Restricting the vocabulary to the top 40K most frequent tokens for each target language introduces no measurable performance loss compared to the previous step, while further improving efficiency. Reducing below 40K works for some languages but does not generalize well across all cases (Appendix E), consistent with Ushio et al. (2023).
4.4 Downstream Performance
Our results show that model compression through knowledge distillation, structured pruning, and vocabulary reduction leads to only a modest performance drop (Table 1). We show that the extent of degradation varies by language and is influenced not just by the amount of distillation data, but also by the strength of the teacher model, which depends on the fine-tuning data size (Appendix F).
Maltese, with the smallest fine-tuning dataset, exhibits the largest performance drops, suggesting weaker knowledge transfer from less-trained teachers and potential reliance on multilingual knowledge. In contrast, Slovak—benefiting from the largest fine-tuning dataset—remains robust even under the maximum compression, indicating that a stronger teacher model facilitates better compression outcomes. Swahili, with a moderate dataset size, follows an intermediate pattern. The results support a relationship between fine-tuning data volume and the model’s resilience to compression.
Adapter capacity: We additionally experiment with varying the reduction factor $r$ to adjust task adapter capacity (see Appendix I and Figure 10). We find that $r=16$ performs well for larger models, while smaller models (e.g., with hidden sizes 564, 456 and 312) benefit from reduced $r$ values (e.g., $r=2$), resulting in modest performance gains. The results for smaller models in Tables 1 and 6 are reported using this lower reduction factor.
5 Related Work
In knowledge distillation, a smaller student model is trained to replicate the behavior of a larger teacher model Hinton et al. (2015), often combining MLM loss with teacher supervision Sun et al. (2019); Sanh et al. (2020). DistilBERT Sanh et al. (2020) reduces model size by selecting every other layer from BERT Devlin et al. (2019) and distills on large corpora using dynamic masking. Patient distillation further improves results by matching intermediate representations Sun et al. (2019).
Recent work has explored distilling multilingual models into compact monolingual models. Singh and Lefever (2022) train student models for languages such as Swahili and Slovenian using a composite loss (distillation, cosine, MLM), and show that distilled models often outperform mBERT while using a reduced vocabulary Abdaoui et al. (2020). Ansell et al. (2023) introduce a two-phase bilingual distillation pipeline, combining general-purpose and task-specific guidance with sparse fine-tuning, outperforming multilingual baselines.
Other studies emphasize the role of initialization. Wibowo et al. (2024) show that copying teacher weights is more effective than random initialization in the context of multilingual distillation, and that MSE outperforms KL divergence for distillation. Cruz (2025) similarly distill mBERT for Tagalog and highlight the nuanced impact of embedding initialization.
6 Conclusion
We present an effective compression pipeline for multilingual encoder models designed for low-resource languages. By integrating staged knowledge distillation, structured pruning, hidden size truncation, and vocabulary reduction, we compress models by up to 92% while maintaining competitive performance - typically within 2–10% of the original - on four downstream tasks.
Limitations and Future Work
Future research should explore optimal strategies for transferring intermediate hidden size knowledge, possibly by applying structured pruning before distillation and identifying the most important components to keep instead of simply truncating them. Additionally, identifying language-specific layers and neurons and leveraging them for more targeted pruning could enhance compression efficiency while maintaining performance. Lastly, more advanced knowledge distillation methods should be experimented with.
Acknowledgements
This research was supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, and by the German Ministry of Education and Research (BMBF) as part of the project TRAILS (01IW24005).
Appendix A KL Divergence vs MSE for Knowledge Distillation
(a) Maltese
(b) Slovak
(c) Swahili
Figure 3: MSE vs. KD validation accuracy for mBERT with the models initialized using the last $k$ layers.
(a) Maltese
(b) Slovak
(c) Swahili
Figure 4: MSE vs. KD validation accuracy for XLM-R with the models initialized using the last $k$ layers.
Appendix B Initialization Strategies for Knowledge Distillation
(a) Maltese
(b) Slovak
(c) Swahili
Figure 5: Validation accuracy for various initialization strategies for mBERT.
(a) Maltese
(b) Slovak
(c) Swahili
Figure 6: Validation accuracy for various initialization strategies for XLM-R.
Appendix C SVD vs. Truncation for Hidden Size Reduction
Figure 7: Validation accuracy comparing SVD and first-$k$ truncation initializations for mBERT models with reduced hidden size ($k$=312). Results show performance with two distillation approaches: using a layer-compressed model (Small teacher) or the original language-adapted model (Big teacher). Student models were trained with identical initialization methods but without knowledge distillation.
Appendix D Alpha Parameter in Knowledge Distillation
(a) mBERT
(b) XLM-R
Figure 8: Validation accuracy curves showing the impact of the alpha parameter on knowledge distillation performance for mBERT and XLM-R on Maltese with the last $k$ and stride initialization strategies for the two models respectively.
We find that the $\alpha$ parameter does not have a significant impact on mBERT during pre-training, with $\alpha=0.5$ yielding consistently good results. For XLM-R, higher values of $\alpha$ (i.e., 0.6 and 0.8), which reduce the strength of the distillation effect, show slightly improved validation accuracy trends compared to lower values. In our experiments, we adopt the default setting of $\alpha=0.5$, leaving a more comprehensive exploration of optimal values across different languages, dataset sizes, and model architectures to future work.
Appendix E Vocabulary Reduction Analysis
(a) Maltese
(b) Slovak
(c) Swahili
Figure 9: Impact of vocabulary reduction on TC performance for mBERT models reduced to a hidden size of 312.
Appendix F Knowledge Distillation Data Sizes
Language KD Data Size (MB) FT Data Size (MB)
Maltese (mt) 238 188
Slovak (sk) 535 1032
Swahili (sw) 402 332
Table 2: Dataset sizes for knowledge distillation (KD) and monolingual fine-tuning (FT) for each language. The language-adapted models are sourced from Gurgurov et al. (2025), and the FT data sizes are as reported by them.
Appendix G Downstream Task Data Sizes
Language Train Validation Test
Text Classification (TC)
Maltese (mt) 701 99 204
Slovak (sk) 701 99 204
Swahili (sw) 701 99 204
Sentiment Analysis (SA)
Maltese (mt) 595 85 171
Slovak (sk) 3560 522 1042
Swahili (sw) 738 185 304
Named Entity Recognition (NER)
Maltese (mt) 100 100 100
Slovak (sk) 20000 10000 10000
Swahili (sw) 1000 1000 1000
Part of Speech Tagging (POS)
Maltese (mt) 1123 433 518
Slovak (sk) 8483 1060 1061
Swahili (sw) 675 134 539
Table 3: Fine-tuning data sizes for each task (Text Classification, Sentiment Analysis, Named Entity Recognition, Part of Speech Tagging) showing train, validation, and test splits across Maltese, Slovak, and Swahili.
Appendix H Downstream Task Hyperparameters
Hyperparameter TC SA NER POS
Learning rate 1e-4 1e-4 3e-4 3e-4
Batch size 16 16 64 64
Epochs 20 20 100 100
Maximum length 256 256 512 512
Table 4: Hyperparameters for task adapter fine-tuning across Text Classification (TC), Sentiment Analysis (SA), and Named Entity Recognition (NER) tasks.
Appendix I Adapter Trainable Parameter Counts
Model Configuration Task Adapter Size
mBERT XLM-R
Base 894,528 894,528
Base-[mt, sk, sw] 894,528 894,528
KD layer red. ×2 447,264 447,264
inter. layer red. → 2048 447,264 447,264
* KD hid. size red. → 564 240,474 240,474
vocab. red. → 40k 240,474 240,474
* KD hid. size red. → 456 156,120 156,120
vocab. red. → 40k 156,120 156,120
* KD hid. size red. → 312 73,122 73,122
vocab. red. → 40k 73,122 73,122
Table 5: Task adapter parameter sizes across different model compression configurations for mBERT and XLM-R with the default reduction factor of 16. When the hidden size is reduced, adapter input/output dimensions decrease proportionally. When the layer count is reduced, fewer adapters are added to the model. All other parameters use the default settings for the Sequential Bottleneck adapter as implemented in AdapterHub.
To examine whether the constrained task adapter capacity, as shown in Table 5, impacts downstream performance in compressed models, we vary the reduction factor $r$, thereby increasing adapter size (see Figure 10). We train task adapters on top of both full adapted models and hidden-size reduced models (564, 456, and 312). For the smallest models (456 and 312), we observe that increasing adapter capacity ($r$=2) leads to improved performance. However, this increase is unnecessary for larger mBERT variants (full and 564), while still beneficial for all small XLM-R models. These results suggest that for smaller models, increasing adapter capacity can yield modest performance gains. Tables 1 and 6 report results using the default reduction rate of 16.
(a) mBERT
(b) XLM-R
Figure 10: Performance of models on TC for Maltese with varying adapter capacity for mBERT and XLM-R.
Appendix J Downstream Results for mBERT
Model Configuration Reduction Size mBERT Avg.
mt sk sw
TC↑ SA↑ NER↑ POS↑ TC↑ SA↑ NER↑ POS↑ TC↑ SA↑ NER↑ POS↑
Base 0% 179M 68.73 65.82 59.96 89.04 85.31 91.98 91.44 97.01 69.60 64.62 83.81 87.61 79.58
Base-[mt, sk, sw] 0% 179M 84.91 73.55 65.00 93.96 86.26 91.87 90.42 96.87 86.68 81.28 82.49 88.69 85.17
KD layer red. ×2 -25% 135M 80.14 73.91 58.96 93.17 85.42 90.38 87.42 96.87 82.78 77.27 80.66 88.42 82.95
inter. layer red. → 2048 -30% 126M 79.03 74.74 58.13 92.74 85.25 90.23 88.48 96.67 83.21 75.90 79.76 88.54 82.72
* KD hid. size red. → 564 -50% 90M 79.47 70.22 61.12 92.64 83.36 90.48 88.07 96.30 83.50 76.10 79.74 88.42 82.45
vocab. red. → 40k -75% 45M 80.17 70.76 61.07 92.46 83.50 90.72 87.73 96.26 84.34 76.01 80.26 88.63 82.66
* KD hid. size red. → 456 -60% 71M 80.23 70.05 57.18 92.13 83.94 90.40 87.49 95.86 85.08 78.57 80.28 88.25 82.45
vocab. red. → 40k -80% 35M 80.95 69.72 55.87 92.00 84.22 90.39 87.38 95.95 83.03 78.50 79.79 88.37 82.18
* KD hid. size red. → 312 -73% 48M 73.14 71.98 39.50 90.33 80.91 90.40 86.48 95.48 81.75 76.48 79.56 87.72 79.48
vocab. red. → 40k -87% 23M 72.98 72.06 40.41 90.18 81.85 90.07 86.19 95.32 81.65 75.97 77.06 87.66 79.28
Table 6: Performance comparison before and after model compression for mBERT across Maltese (mt), Slovak (sk), and Swahili (sw). Task metrics: TC = Text Classification, SA = Sentiment Analysis, NER = Named Entity Recognition, POS = Part-of-Speech Tagging. All scores are averaged over three independent runs with different random seeds.