### v2

# On Multilingual Encoder Language Model Compression for Low-Resource Languages

Abstract
In this paper, we combine two-step knowledge distillation, structured pruning, truncation, and vocabulary trimming for extremely compressing multilingual encoder-only language models for low-resource languages. Our novel approach systematically combines existing techniques and takes them to the extreme, reducing layer depth, feed-forward hidden size, and intermediate layer embedding size to create significantly smaller monolingual models while retaining essential language-specific knowledge. We achieve compression rates of up to 92% while maintaining competitive performance, with average drops of 2–10% for moderate compression and 8–13% at maximum compression in four downstream tasks, including sentiment analysis, topic classification, named entity recognition, and part-of-speech tagging, across three low-resource languages. Notably, the performance degradation correlates with the amount of language-specific data in the teacher model, with larger datasets resulting in smaller performance losses. Additionally, we conduct ablation studies to identify the best practices for multilingual model compression using these techniques.
1 Introduction
Small multilingual encoder language models (LMs), such as mBERT devlin-etal-2019-bert, XLM-R conneau-etal-2020-unsupervised, and Glot-500m imanigooghari-etal-2023-glot500, have demonstrated strong performance across a diverse range of low-resource languages pmlr-v119-hu20b; asai-etal-2024-buffet, often outperforming large-scale proprietary models on various sequential tasks adelani-etal-2024-sib; gurgurov2025smallmodelsbigimpact. However, even these relatively compact multilingual models may still be excessively large for use in individual languages due to redundant capacity and expensive inference singh-lefever-2022-student; cruz-2025-extracting.
Figure 1: Overview of our multilingual model compression methodology. We use (1) knowledge distillation to reduce layers, (2) structured pruning to eliminate redundant feed-forward network width, and (3) hidden size reduction and another round of knowledge distillation from the previous student model. Finally, (4) vocabulary trimming is applied to retain language-specific tokens.
To address this, we propose a novel combination of model compression approaches for transforming multilingual encoder-only models into maximally small, efficient, language-specific alternatives while retaining competitive performance. Our methodology integrates knowledge distillation hinton2015distillingknowledgeneuralnetwork, structured pruning kim-hassan-2020-fastformers; hou2020dynabert, weight truncation, and vocabulary trimming abdaoui-etal-2020-load; ushio-etal-2023-efficient to systematically reduce model size by compressing the depth (number of layers), feed-forward intermediate width, hidden size, and tokenizer vocabulary. Our experiments demonstrate that this pipeline achieves compression rates of up to 92%, with performance drops of 2-10% for moderate compression (up to 87%) and 8-13% at maximum compression on downstream tasks such as sentiment analysis, topic classification, named entity recognition, and part-of-speech tagging. Notably, for moderate compression levels, the extent of degradation depends more on the strength of the teacher model than on the compression itself.
Beyond compression, we investigate the impact of using multilingual versus monolingual teacher models, evaluate different initialization strategies for knowledge distillation, and analyze additional compression variables. Our findings contribute to the development of highly efficient, environmentally friendly models Strubell_Ganesh_McCallum_2020 for low-resource languages and explore how strongly models can be compressed. The code for our experiments is made publicly available at https://github.com/d-gurgurov/Multilingual-LM-Disitillation.
2 Methodology
In this section, we present our multilingual model compression strategy, illustrated in Figure 1. Our approach combines several existing compression techniques in a novel way that, to the best of our knowledge, has not been explored in this combination within the multilingual context.
2.1 Layer Reduction via Knowledge Distillation
We reduce the number of transformer layers in the teacher model by half to obtain an initial compact student model sanh2020distilbertdistilledversionbert. The student is initialized with the layers of the teacher and trained using a combination of Masked Language Modeling (MLM) devlin-etal-2019-bert and Mean Squared Error (MSE) loss for knowledge distillation hinton2015distillingknowledgeneuralnetwork for 10 epochs. Both losses are weighted equally ($\alpha$=0.5, though other values were explored; see Appendix 8). The teacher is a multilingual encoder fine-tuned on the target language (see Section 4).
2.2 Width Reduction via Structured Pruning
We apply structured pruning kim-hassan-2020-fastformers to reduce the intermediate size of the feed-forward layers from 3072 to 2048. Neuron importance is estimated using first-order gradient information accumulated from forward and backward passes over MLM validation data. At each layer, neurons are ranked by their absolute gradient values, and the least important ones are removed based on a target pruning ratio. The remaining neurons are then reordered to preserve model functionality. For consistency, the same pruning ratio is applied across all layers.
2.3 Hidden Size Compression with Secondary Knowledge Distillation
We compress the hidden embedding dimension from 768 to either 312, 456, or 564 via truncation, retaining the first $k$ dimensions. [^1] The hidden size must be divisible by the number of attention heads. A second round of knowledge distillation is then performed, using the width-reduced model from the previous step as the new teacher, similar to 10254426, with training for 10 epochs.
2.4 Vocabulary Reduction
We reduce the vocabulary size by selecting the top 40,000 most frequent tokens from a target-language corpus, along with their corresponding embeddings ushio-etal-2023-efficient. This ensures that the resulting model retains only language-specific tokens, which significantly reduces the overall model size.
3 Experiments
Below, we describe the datasets, languages, tasks, and baseline systems used in our evaluation.
3.1 Knowledge Distillation Data
We use GlotCC kargaran2025glotccopenbroadcoveragecommoncrawl, a large-scale multilingual corpus derived mainly from CommonCrawl wenzek-etal-2020-ccnet, as the primary dataset for both stages of knowledge distillation. Data distributions for the selected languages are reported in Appendix F. We use GlotCC for training, and the FLORES-200 development set nllbteam2022languageleftbehindscaling for validation during training.
3.2 Languages and Tasks
We evaluate our models on four tasks: Topic Classification (TC), Sentiment Analysis (SA), Named Entity Recognition (NER), and Part-of-Speech Tagging (POS), covering three low-resource languages–Maltese, Slovak, and Swahili joshi2020state. For TC, we use the 7-class SIB-200 dataset adelani-etal-2024-sib, and for SA, we compile binary sentiment datasets from multiple sources dingli2016sentiment; cortis-davis-2019-social; pecar-etal-2019-improving; Muhammad2023AfriSentiAT; muhammad2023semeval. For NER, we use WikiANN pan-etal-2017-cross, and for POS, we use Universal Dependencies v2.15 de-marneffe-etal-2021-universal and MasakhaPOS dione-etal-2023-masakhapos. For all tasks, we train Sequential Bottleneck task adapters pfeiffer-etal-2020-mad with fixed hyperparameters (see Appendix H). Performance is measured using macro-averaged F1 sokolova2006beyond for TC and SA, and "seqeval" F1 seqeval for NER and POS.
3.3 Models and Baselines
We compress two encoder multilingual models–mBERT devlin-etal-2019-bert and XLM-R-base conneau-etal-2020-unsupervised–adapted to target languages through fine-tuning on language-specific data, and compare the reduced models to two baselines: (1) the original, non-adapted models, and (2) language-adapted versions. In both cases, we train an identical task adapter using the same task-specific datasets as for the compressed models.
4 Findings
Our key findings are outlined below.
4.1 Distillation
Distilling knowledge from a multilingual teacher into a monolingual student model is less effective than using a target-language adapted teacher, as evidenced by the differences in validation accuracies shown in Figure 2. This discrepancy possibly stems from the multilingual teacher’s broad cross-lingual representations, which are not directly aligned with the requirements of a monolingual student. In contrast, monolingual teachers provide more targeted, language-specific representations, resulting in better student performance.
(a) Maltese (mBERT)
(b) Maltese (XLM-R)
(c) Slovak (mBERT)
(d) Slovak (XLM-R)
(e) Swahili (mBERT)
(f) Swahili (XLM-R)
Figure 2: First-step KD validation accuracies for mBERT and XLM-R with models initialized using the last $k$ layers. mBERT- and XLM-R-mt, sk, sw refer to models adapted to the target language; distilled denotes models trained with distillation loss, while student refers to identically trained models without distillation loss. The best accuracy is in all cases achieved when distilling from a target-language adapted model.
Distillation loss: We compare KL divergence and MSE as distillation loss functions, and observe that MSE leads to better and faster convergence (Appendix A), in line with prior work kim2021comparingkullbackleiblerdivergencemean; nityasya2022studentbestcomprehensiveknowledge.
4.2 Weight Initialization
Weight initialization plays a crucial role in training the student model, with knowledge distillation providing only a marginal additional performance improvement (Figure 2). This partly aligns with the findings of wibowo2024privilegedstudentsvalueinitialization, who explored distilling multilingual abilities for multilingual tasks, whereas our focus is on monolingual distillation. Training a student-sized model initialized with teacher weights, but without knowledge distillation, results in a slight performance drop compared to a fully distilled model.
Initialization Strategies: Among various initialization strategies, initializing the student with the last $k$ layers for mBERT and every other layer (stride) for XLM-R consistently outperforms alternatives such as using the first $k$ layers and combining first and last layers (Appendix B). Random initialization performs significantly worse, emphasizing the importance of weight reuse sun2019patientknowledgedistillationbert; singh-lefever-2022-student.
Compression Stage Params Size Task Performance (F1) Avg
Maltese Slovak Swahili
TC SA NER POS TC SA NER POS TC SA NER POS
Baselines
   Multilingual 279M 279M 68.1 56.0 54.3 89.9 88.1 95.6 91.1 97.3 78.4 81.5 84.6 89.4 81.2
[…]
Compression Pipeline (minimal degradation)
   Layer reduction 236M (-15%) 236M 84.0 77.2 63.5 94.3 86.3 92.9 90.1 96.3 82.9 81.3 82.9 89.2 85.1
   + FFN pruning 226M (-20%) 226M 84.7 78.6 60.1 94.2 86.1 93.4 90.0 96.1 82.4 82.7 83.6 89.5 85.1
   + Hidden 564 163M (-40%) 163M 83.4 74.9 53.0 93.7 84.9 92.7 89.1 96.8 85.8 81.0 80.8 89.4 83.8
[…]
Further compression (moderate degradation)
   + Hidden 456 131M (-53%) 131M 78.5 69.9 62.5 92.7 86.0 93.0 88.3 96.3 83.1 79.3 80.7 88.9 83.3
   + Vocabulary 35M (-87%) 35M 78.5 70.7 63.3 92.5 86.1 92.9 88.4 96.3 82.5 79.0 80.2 89.0 83.3
Maximum compression (higher degradation)
[…]
Table 1: Progressive compression of XLM-R-base. Stages are grouped by degradation level. Highlighted rows indicate the baseline (gray) and optimal compression point (green, 85% reduction with 2.5% drop). Maximum compression rows (red) show higher degradation rates (7.6% drop). All F1 scores are averaged over 3 independent runs with different random seeds mBERT in Appendix J.
4.3 Pruning and Truncation
Distilled models can be compressed further using structured pruning, hidden size reduction, and vocabulary trimming, while maintaining competitive performance.
Intermediate size reduction: Reducing the intermediate size of feed-forward layers from 3072 to 2048 via structured pruning results in negligible performance loss (Table 1). However, more aggressive reductions degrade quality significantly, making 2048 a practical lower bound. We do not prune attention heads, as removing even a minimal number (e.g., three) causes severe degradation (>50% performance drop in preliminary experiments).
Hidden size reduction: We reduce the hidden embedding size to 564, 456, and 312, truncating it to the first $k$ dimensions. Training is performed under the supervision of the student from the previous stage. We find that using the original teacher leads to worse results, possibly due to the bigger knowledge gap 10254426. We also tested SVD-based dimensionality reduction but found truncation to be more effective (see Appendix C).
Vocabulary trimming: Restricting the vocabulary to the top 40K most frequent tokens for each target language introduces no measurable performance loss compared to the previous step, while further improving efficiency. Reducing below 40K works for some languages but does not generalize well across all cases (Appendix E), consistent with ushio-etal-2023-efficient.
4.4 Downstream Performance
Our results show that model compression through knowledge distillation, structured pruning, and vocabulary reduction leads to modest performance drops (Tables 1 and 6). Below, we report results for XLM-R; results for mBERT follow similar patterns and are presented in Appendix J.
Language-specific resilience: The extent of degradation varies by language and correlates with teacher model quality. At maximum compression (92% parameter reduction), Slovak (1032MB fine-tuning (FT) data) experiences only a 2.9% performance drop, Swahili (332MB) shows a 5.2% drop, while Maltese (188MB) degrades by 19.2%. This pattern demonstrates that stronger teacher models–trained on larger datasets–enable more robust compression outcomes.
Task-specific patterns: Different tasks exhibit varying compression sensitivities. POS tagging shows the highest resilience across all languages, with performance drops of only 4-13% at 92% compression. Conversely, NER demonstrates steeper degradation, particularly for Maltese (69.2 → 37.1 F1). This severe drop is likely compounded by the extremely small Maltese NER training set (100 examples vs. 20,000 for Slovak), indicating that sequence labeling tasks are especially vulnerable to compression in low-resource settings. In contrast, sentence-level classification tasks such as SA and TC remain relatively stable under heavy compression, with performance decreases below 10% even at 85–90% size reduction.
Optimal compression trade-offs: The 85% compression level (hidden size 564 with 40k vocabulary) offers the best balance for most scenarios, with only a 2.5% average performance drop (84.3 vs 86.5 avg F1). For high-resource languages like Slovak, even 87% compression incurs only a 3.8% drop. Notably, vocabulary trimming often yields slight improvements (e.g., Maltese TC: 84.11 vs 83.43 F1), suggesting it reduces vocabulary noise while compensating for hidden size reduction.
Staged compression effects: Layer reduction (15%) and intermediate size pruning (20%) induce minimal degradation (<2% drop), with the primary performance impact occurring during hidden size reduction. Performance degrades gradually up to 85% compression, but deteriorates more rapidly beyond this threshold (4-6% drop per additional stage).
Adapter capacity: We experiment with varying the reduction factor $r$ to adjust task adapter capacity (Appendix I, Figure 10). While $r=16$ suffices for larger models, smaller models (hidden sizes 564, 456, 312) benefit from lower $r$ values ($r=2$), yielding modest performance gains. Results in Tables 1 and 6 use $r=2$ for these compressed models.
5 Related Work
In knowledge distillation, a smaller student model is trained to replicate the behavior of a larger teacher model hinton2015distillingknowledgeneuralnetwork, often combining MLM loss with teacher supervision sun2019patientknowledgedistillationbert; sanh2020distilbertdistilledversionbert. DistilBERT sanh2020distilbertdistilledversionbert reduces model size by selecting every other layer from BERT devlin-etal-2019-bert and distills on large corpora using dynamic masking. Patient distillation further improves results by matching intermediate representations sun2019patientknowledgedistillationbert.
Recent work has explored distilling multilingual models into compact monolingual models. singh-lefever-2022-student train student models for languages such as Swahili and Slovenian using a composite loss (distillation, cosine, MLM), and show that distilled models often outperform mBERT while using a reduced vocabulary abdaoui-etal-2020-load. ansell2023distillingefficientlanguagespecificmodels introduce a two-phase bilingual distillation pipeline, combining general-purpose and task-specific guidance with sparse fine-tuning, outperforming multilingual baselines.
Other studies emphasize the role of initialization. wibowo2024privilegedstudentsvalueinitialization show that copying teacher weights is more effective than random initialization in the context of multilingual distillation, and that MSE outperforms KL divergence for distillation. cruz-2025-extracting similarly distill mBERT for Tagalog and highlight the nuanced impact of embedding initialization.
6 Conclusion
We present an effective compression pipeline for multilingual encoder models designed for low-resource languages. By integrating staged knowledge distillation, structured pruning, hidden size truncation, and vocabulary reduction, we compress models by up to 92% while maintaining competitive performance, typically within 2–10% of the original for moderate compression and 8–13% at maximum compression, on four downstream tasks.
Limitations
Our evaluation is limited to three low-resource languages and four downstream tasks, which may affect generalizability to other languages and task types. The compression pipeline requires target-language data for teacher adaptation, making it less suitable for truly low-resource languages with minimal corpora. We focus exclusively on encoder-only models (mBERT and XLM-R), and our structured pruning only targets feed-forward layers, leaving attention head pruning unexplored due to performance degradation.
Acknowledgments
This research was supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, by the German Ministry of Education and Research (BMBF) as part of the project TRAILS (01IW24005), and by lorAI - Low Resource Artificial Intelligence, a project funded by the European Union under GA No.101136646.
Appendix A KL Divergence vs MSE for Knowledge Distillation
(a) Maltese
(b) Slovak
(c) Swahili
Figure 3: MSE vs. KD validation accuracy for mBERT with the models initialized using the last $k$ layers.
(a) Maltese
(b) Slovak
(c) Swahili
Figure 4: MSE vs. KD validation accuracy for XLM-R with the models initialized using the last $k$ layers.
Appendix B Initialization Strategies for Knowledge Distillation
(a) Maltese
(b) Slovak
(c) Swahili
Figure 5: Validation accuracy for various initialization strategies for mBERT.
(a) Maltese
(b) Slovak
(c) Swahili
Figure 6: Validation accuracy for various initialization strategies for XLM-R.
Appendix C SVD vs. Truncation for Hidden Size Reduction
Figure 7: Validation accuracy comparing SVD vs. first-$k$ truncation for hidden size reduction to 312. “Small teacher” refers to the layer-compressed (6-layer) model; “Big teacher” is the original 12-layer language-adapted model. Truncation consistently outperforms SVD regardless of teacher size.
Appendix D Alpha Parameter in Knowledge Distillation
(a) mBERT
(b) XLM-R
Figure 8: Validation accuracy curves showing the impact of the alpha parameter on knowledge distillation performance for mBERT and XLM-R on Maltese with the last $k$ and stride initialization strategies for the two models respectively.
We find that the $\alpha$ parameter does not have a significant impact on mBERT during pre-training, with $\alpha=0.5$ yielding consistently good results. For XLM-R, higher values of $\alpha$ (i.e., 0.6 and 0.8), which reduce the strength of the distillation effect, show slightly improved validation accuracy trends compared to lower values. In our experiments, we adopt the default setting of $\alpha=0.5$, leaving a more comprehensive exploration of optimal values across different languages, dataset sizes, and model architectures to future work.
Appendix E Vocabulary Reduction Analysis
(a) Maltese
(b) Slovak
(c) Swahili
Figure 9: Impact of vocabulary reduction on TC performance for mBERT models reduced to a hidden size of 312.
Appendix F Knowledge Distillation Data Sizes
Language KD Data Size (MB) FT Data Size (MB)
Maltese (mt) 238 188
Slovak (sk) 535 1032
Swahili (sw) 402 332
Table 2: Dataset sizes for knowledge distillation (KD) and monolingual fine-tuning (FT) for each language. The language-adapted models are sourced from gurgurov2025smallmodelsbigimpact, and the FT data sizes are as reported by them.
Appendix G Downstream Task Data Sizes
Language Train Validation Test
Text Classification (TC)
Maltese (mt) 701 99 204
Slovak (sk) 701 99 204
Swahili (sw) 701 99 204
Sentiment Analysis (SA)
Maltese (mt) 595 85 171
Slovak (sk) 3560 522 1042
Swahili (sw) 738 185 304
Named Entity Recognition (NER)
Maltese (mt) 100 100 100
Slovak (sk) 20000 10000 10000
Swahili (sw) 1000 1000 1000
Part of Speech Tagging (POS)
Maltese (mt) 1123 433 518
Slovak (sk) 8483 1060 1061
Swahili (sw) 675 134 539
Table 3: Fine-tuning data sizes for each task (Text Classification, Sentiment Analysis, Named Entity Recognition, Part of Speech Tagging) showing train, validation, and test splits across Maltese, Slovak, and Swahili.
Appendix H Downstream Task Hyperparameters
Hyperparameter TC SA NER POS
Learning rate 1e-4 1e-4 3e-4 3e-4
Batch size 16 16 64 64
Epochs 20 20 100 100
Maximum length 256 256 512 512
Table 4: Hyperparameters for task adapter fine-tuning across Text Classification (TC), Sentiment Analysis (SA), and Named Entity Recognition (NER) tasks.
Appendix I Adapter Trainable Parameter Counts
Model Configuration Task Adapter Size
mBERT XLM-R
Base 894,528 894,528
Base-[mt, sk, sw] 894,528 894,528
KD layer red. ×2 447,264 447,264
inter. layer red. → 2048 447,264 447,264
* KD hid. size red. → 564 240,474 240,474
vocab. red. → 40k 240,474 240,474
* KD hid. size red. → 456 156,120 156,120
vocab. red. → 40k 156,120 156,120
* KD hid. size red. → 312 73,122 73,122
vocab. red. → 40k 73,122 73,122
Table 5: Task adapter parameter sizes across different model compression configurations for mBERT and XLM-R with the default reduction factor of 16. When the hidden size is reduced, adapter input/output dimensions decrease proportionally. When the layer count is reduced, fewer adapters are added to the model. All other parameters use the default settings for the Sequential Bottleneck adapter as implemented in AdapterHub.
To examine whether the constrained task adapter capacity, as shown in Table 5, impacts downstream performance in compressed models, we vary the reduction factor $r$, thereby increasing adapter size (see Figure 10). We train task adapters on top of both full adapted models and hidden-size reduced models (564, 456, and 312). For the smallest models (456 and 312), we observe that increasing adapter capacity (rr=2) leads to improved performance. However, this increase is unnecessary for larger mBERT variants (full and 564), while still beneficial for all small XLM-R models. These results suggest that for smaller models, increasing adapter capacity can yield modest performance gains. Tables 1 and 6 report results using the default reduction rate of 16.
(a) mBERT
(b) XLM-R
Figure 10: Performance of models on TC for Maltese with varying adapter capacity for mBERT and XLM-R.
Appendix J Downstream Results for mBERT
Compression Stage Params Size Task Performance (F1) Avg
Maltese Slovak Swahili
TC SA NER POS TC SA NER POS TC SA NER POS
Baselines
   Multilingual 179M 179M 68.7 65.8 60.0 89.0 85.3 92.0 91.4 97.0 69.6 64.6 83.8 87.6 79.6
[…]
Compression Pipeline (minimal degradation)
   Layer reduction 135M (-25%) 135M 80.1 73.9 59.0 93.2 85.4 90.4 87.4 96.9 82.8 77.3 80.7 88.4 83.0
   + FFN pruning 126M (-30%) 126M 79.0 74.7 58.1 92.7 85.3 90.2 88.5 96.7 83.2 75.9 79.8 88.5 82.7
   + Hidden 564 90M (-50%) 90M 79.5 70.2 61.1 92.6 83.4 90.5 88.1 96.3 83.5 76.1 79.7 88.4 82.5
[…]
Further compression (moderate degradation)
   + Hidden 456 71M (-60%) 71M 80.2 70.1 57.2 92.1 83.9 90.4 87.5 95.9 85.1 78.6 80.3 88.3 82.5
   + Vocabulary 35M (-80%) 35M 81.0 69.7 55.9 92.0 84.2 90.4 87.4 96.0 83.0 78.5 79.8 88.4 82.2
Maximum compression (higher degradation)
[…]
Table 6: Progressive compression of mBERT. Stages are grouped by degradation level. Highlighted rows indicate the baseline (gray) and optimal compression point (green, 75% reduction with 2.5% drop). Maximum compression rows (red) show significant degradation (5.9% drop). TC=Topic Classification, SA=Sentiment Analysis, NER=Named Entity Recognition, POS=Part-of-Speech Tagging. F1 scores averaged over 3 runs.